# Infinite systems of equations in abelian and nilpotent groups111 This work was supported by the Russian Science Foundation, project no. 22-11-00075

### v1

@@ 1 Introduction
In this work $\mathbb{Z}_{p}$ for a prime number $p$ denotes the field of order $p$.
@@
Let $\pi$ be some set of prime numbers (maybe not of all prime numbers). A finite system of equations is called $\pi$-nonsingular if it is $p$-nonsingular for every prime number $p\in\pi$. By definition $\varnothing$-nonsingular system is just a nonsingular system.
@@
In particular, a single equation $w(x)=1$ is a single variable is nonsingular ($p$-nonsingular, unimodular respectively) if the exponent sum of $x$ in $w=1$ is not equal to zero (is not divisible by $p$, is equal to ±$\pm 1$ respectively).
@@
If a system of equations over a group is $p$-nonsingular for some prime number $p$ then it is also nonsingular, as having a nontrivial combination of integer rows over $\mathbb{Z}$, which is equal to zero, implies that there is a likewise combination in which at least one coefficient is not divisible by $p$. Particularly, unimodular systems of equations are nonsingular. Nonsingularity of a finite system of equations can be interpreted as existence of a non-zero minor of rank $k$ in the matrix of exponent sums of variables, where $k$ is the number of equations. Similarly, $p$-nonsingularity can be seen as existence of a likewise minor which is not divisible by $p$. Hence, if a system of equations is finite, then its nonsingularity implies its $p$-nonsingularity for some prime number $p$. This, however, is not true for infinite systems: the system $\{x_{p}^{p}=1\},$ where $p$ ranges over all prime numbers, is nonsingular but is $p$-singular for every prime number $p$.
@@
  • $p$-nonsingular systems are solvable over locally $p$-indicable groups [9], see also [19]. Locally $p$-indicable group is a group whose every nontrivial finitely generated subgroup admits an epimorphism onto the additive group of $\mathbb{Z}_{p}$;
@@
In fact, the solvability of a system $\{w_{i}=1\}_{i\in I}$ of equations in variables $\{x_{j}\}_{j\in J}$ over a group $G$ is equivalent to the injectivity of a natural homomorphism
@@
Consider a finite $\pi$-nonsingular system of equations over a group and the matrix of the exponent sums of variables in these equations. Suppose that there are $k$ equations and $n$ variables in the system and $k<n$ (i.e. the matrix of the system is not square). Using transformations (i.e. multiplying one equation by another or by an inverse of another and rearranging equations) of equations in the system (which lead to transformations of rows in the matrix) and changes of variables (which lead to transformations of columns in the matrix), we can transform the matrix to a triangular matrix in which the first $k$ columns modulo $p$ are linearly independent over $\mathbb{Z}_{p}$ for each $p\in\pi$. Therefore this transformed system of equations remains $\pi$-nonsingular even after elimination of all variables except the first $k$ ones (for example, the other variables can be changed to $1$). Thus, any finite $\pi$-nonsingular system of equations can be reduced to a likewise system in which there are as many equations as there are variables (i.e. the matrix of the system is square). So it can be assumed for convenience that finite nonsingular, $\pi$-nonsingular or unimodular systems are ‘‘square’’.
@@
If a group is finite, then every finite nonsingular system of equations over this group is solvable in a finite group as well [10]. If a group is abelian, then every finite nonsingular system of equations over this group is solvable in an abelian group, namely in the divisible hull of the group. For example, one can solve the system by Cramer’s rule (having in mind that it is possible to divide in a divisible abelian group, not necessarily uniquely). Moreover, if a finite system of equations over an abelian groups is unimodular, then the system is solvable in this abelian group itself. This is also true for all nilpotent group: it is a special case of Shmel’kin’s theorem. Before we formulate this theorem, let us remind a couple of definitions.
@@
Particularly, a $\varnothing$-divisible group is just any group.
@@
Particularly, a $\varnothing$-torsion-free group is just any group.
@@
It follows that non-square finite $\pi^{\prime}$-nonsingular systems have a solution (although not unique) in $G$ as well, as discussed above.
@@
Also, in this section we prove the same generalization of Shmel’kin’s theorem for divisible nilpotent group (not even necessarily torsion free).
@@
The author thanks the Theoretical Physics and Mathematics Advancement Foundation ‘‘BASIS’’. The author also thanks Anton Klyachko for valuable remarks and the mentorship.
@@ 2.1 Negative case
First consider the case where $I$ is finite. Then some of $A_{p}$ has unbounded period (otherwise the period of $A$ is bounded). If any unimodular system of equations over $A$ has a solution in $A$, then this is true for any quotient group of $A$. However, Proposition 1 shows that for $A_{p}$ this fact is not true. So, not all unimodular systems of equations over $A$ are solvable in $A$. One can even give an example of a system which has no solutions in $A$: if the projection of the system onto $A_{p}$ is the system from the proof of Proposition 1, then the system itself is unsolvable in $A$.
@@ 2.2 Positive case
Any $p$-nonsingular system of equations over the abelian group $A$ of prime period $p$ has a solution in $A$ itself.
@@
Look at the quotient group $A/pA$. Take the induced system $\{m_{i}=\overline{a}_{i}\}_{i\in I}$ of equations over $A/pA$, where $\overline{a}_{i}$ is the image of $a_{i}$ under taking the quotient $A\to A/pA$. It is a $p$-nonsingular system of equations over the abelian group $A/pA$ of period $p$, therefore by Proposition 2 the system has a solution $\{\bar{c}_{j}\}_{j\in J}$ in $A/pA$.
@@
where $A_{p_{i}}$ is the $p_{i}$-component of $A$ (which has bounded period $p_{i}^{n_{i}}$). For the components $A_{p_{i}}$ the statement of the lemma follows from Proposition 3, therefore it is also true for their direct sum. ∎
@@ 2.3 The final criterion
Nowe we can prove the criterion formulated in the introduction.
@@
Decompose $A$ into a direct sum of the divisible abelian group and the reduced one. The divisible abelian group has a solution of every nonsingular system of equations over itself (by Lemma 1). Therefore, nonsingular (particularly, unimodular) systems of equations have a solution in $A$ if and only if the same is true for its reduced part. Hence we can assume $A$ to be reduced.
@@ 3 Nilpotent groups
Now we switch back to the multiplicative notation. The obtained results can be used to research solvability of systems of equations in nilpotent groups.
@@ 3.1 Periodic nilpotent groups
Take a look at the system $\{\overline{w}_{i}=1\}_{i\in I}$ over $G/Z(G)$, where for each $i\in I$ the word $\overline{w}_{i}$ is the word $w_{i}$ with coefficients changed to their images under taking the quotient $G\to G/Z(G)$. The system $\{\overline{w}_{i}=1\}$ is a unimodular system of equations over $G/Z(G)$ (which is a nilpotent group of class $l$ and of bounded period), therefore the system is solvable in the quotient group $G/Z(G)$ itself by induction hypothesis.
This means that we can make a substitution of the variables in the system $\{w_{i}=1\}$ over $G$ such that for each $i\in I$ product $b_{i}$ of coefficients of the word $w_{i}$ lies in $Z(G)$ (i.e. the image of $b_{i}$ under taking the quotient $G\to G/Z(G)$ is $1$). So we can assume that the words $w_{i}$ initially have this property (that is, the products of their coefficients lie in $Z(G)$).
@@ 3.2 Divisible nilpotent groups
Take a look at the system $\{\overline{w}_{i}=1\}_{i\in I}$ over $G/Z(G)$, where for each $i\in I$ the word $\overline{w}_{i}$ is the word $w_{i}$ with coefficients changed to their images under taking the quotient $G\to G/Z(G)$. The system $\{\overline{w}_{i}=1\}$ is a nonsingular system of equations over $G/Z(G)$ (which is a divisible nilpotent group of class $l$), therefore the system is solvable in the quotient group $G/Z(G)$ itself by induction hypothesis.
This means that we can make a substitution of the variables in the system $\{w_{i}=1\}$ over $G$ such that for each $i\in I$ the product $b_{i}$ of coefficients of the word $w_{i}$ lies in $Z(G)$ (i.e. the image of $b_{i}$ under taking the quotient $G\to G/Z(G)$ is $1$). So we can assume that the words $w_{i}$ initially have this property (that is the products of their coefficients lie in $Z(G)$).
@@ 4 Open questions
The same question for mixed abelian groups.

### v3

@@ 1 Introduction
In this work $\mathbb{Z}_{p}$ for a prime number $p$ denotes the field of order $p$. Here we refer to the exponent of a group $G$ as the period of $G$.
@@
Let $\pi$ be some set of prime numbers (maybe not of all prime numbers). A finite nonsingular system of equations is called $\pi$-nonsingular if it is $p$-nonsingular for every prime number $p\in\pi$. Patricularly, $\varnothing$-nonsingularity of a system is the same as its nonsingularity.
@@
In particular, a single equation $w(x)=1$ in a single variable is nonsingular ($p$-nonsingular, unimodular respectively) if the exponent sum of $x$ in $w=1$ is not equal to zero (is not divisible by $p$, is equal to ±$\pm 1$ respectively).
@@
If a system of equations over a group is $p$-nonsingular for some prime number $p$ then it is also nonsingular, as having a nontrivial combination of integer rows over $\mathbb{Z}$ which is equal to zero implies that there is a likewise combination in which at least one coefficient is not divisible by $p$. Particularly, unimodular systems of equations are nonsingular. Nonsingularity of a finite system of equations can be interpreted as existence of a non-zero minor of rank $k$ in the matrix of exponent sums of variables, where $k$ is the number of equations. Similarly, $p$-nonsingularity can be seen as existence of a likewise minor which is not divisible by $p$. Hence, if a system of equations is finite, then its nonsingularity implies its $p$-nonsingularity for some prime number $p$. This, however, is not true for infinite systems: the system $\{x_{p}^{p}=1\},$ where $p$ ranges over all prime numbers, is nonsingular but is $p$-singular for every prime number $p$.
@@
  • $p$-nonsingular systems are solvable over locally $p$-indicable groups [9], see also [19]. A locally $p$-indicable group is a group whose every nontrivial finitely generated subgroup admits an epimorphism onto the additive group of $\mathbb{Z}_{p}$;
@@
In fact, the solvability of a system $\{w_{i}=1\}_{i\in I}$ of equations in variables $\{x_{j}\}_{j\in J}$ over a group $G$ is equivalent to the injectivity of the natural homomorphism
@@
Consider a finite $\pi$-nonsingular system of equations over a group and the matrix of the exponent sums of variables in these equations. Suppose that there are $k$ equations and $n$ variables in the system and $k<n$ (i.e. the matrix of the system is not square). Using transformations (i.e. multiplying one equation by another or by an inverse of another and rearranging equations) of equations in the system (which lead to transformations of rows in the matrix) and changes of variables (which lead to transformations of columns in the matrix), we can transform the matrix to a triangular matrix in which the first $k$ columns modulo $p$ are linearly independent over $\mathbb{Z}_{p}$ for each $p\in\pi$. Therefore, this transformed system of equations remains $\pi$-nonsingular even after elimination of all variables except the first $k$ ones (for example, the other variables can be changed to $1$). Thus, any finite $\pi$-nonsingular system of equations can be reduced to a likewise system in which there are as many equations as there are variables (i.e. the matrix of the system is square). So it can be assumed for convenience that finite nonsingular, $\pi$-nonsingular or unimodular systems are ‘‘square’’.
@@
If a group is finite, then every finite nonsingular system of equations over this group is solvable in a finite group as well [10]. If a group is abelian, then every finite nonsingular system of equations over this group is solvable in an abelian group, namely in the divisible hull of the group. For example, one can solve the system by Cramer’s rule (having in mind that it is possible to divide in a divisible abelian group, not necessarily uniquely). Moreover, if a finite system of equations over an abelian group is unimodular, then the system is solvable in this abelian group itself. This is also true for all nilpotent groups: it is a special case of Shmel’kin’s theorem. Before we formulate this theorem, let us remind a couple of definitions.
@@
Particularly, any group is $\varnothing$-divisible.
@@
Particularly, any group is $\varnothing$-torsion free.
@@
It follows that every non-square finite $\pi^{\prime}$-nonsingular system of equations over $G$ has a solution (although not unique) in $G$ as well, as discussed above.
@@
Indeed, suppose that the said group contains a solution $\{\widetilde{x},\widetilde{y}_{2},\widetilde{y}_{3},\widetilde{y}_{5},\ldots\}$ of this system. Then, as $\widetilde{x}+2\widetilde{y}_{2}=(1,0,0,\ldots)$ and the first coordinate of $2\widetilde{y}_{2}$, which is an element of $\mathbb{Z}_{2}$ multiplied by $2$, equals 00, we get that the first coordinate of $\widetilde{x}$ is not equal to 00. Also, since $\widetilde{x}+3\widetilde{y}_{3}=(0,1,0,\ldots)$ and the second coordinate of $3\widetilde{y}_{3}$ is equal to 00, as it is an element of $\mathbb{Z}_{3}$ multiplied by $3$, the second coordinate of $\widetilde{x}$ is not equal to 00 as well. Likewise we get that all the coordinates of $\widetilde{x}$ are not equal to 00. In other words, $\widetilde{x}$ has an infinite number of coordinates not equal to 00 and hence can not belong to $\mathbb{Z}_{2}\oplus\mathbb{Z}_{3}\oplus\mathbb{Z}_{5}\oplus\ldots$. Thus we arrive at a contradiction, so the group $\mathbb{Z}_{2}\oplus\mathbb{Z}_{3}\oplus\mathbb{Z}_{5}\oplus\ldots$ does not have a solution to the system above.
@@
Here the reduced part of an abelian group $A$ is any group $C$ such that $A=C\oplus D$ with $D$ divisible and $C$ reduced. Such decomposition always exists, and in every such decomposition $D$ is the largest divisible subgroup of $A$ and $C$ is isomorphic to $A/D$, so the reduced part of an abelian group is unique up to isomorphism.
@@
Also, in this section we prove the same generalization of Shmel’kin’s theorem for divisible nilpotent groups (not even necessarily torsion free).
@@
The author thanks the Theoretical Physics and Mathematics Advancement Foundation ‘‘BASIS’’. The author also thanks Anton Klyachko for valuable remarks, the mentorship and for the example from Question 1. The author thanks an anonymous referee for useful comments.
@@ 2.1 Negative case
First consider the case where $I$ is finite. Then some component $A_{p}$ has unbounded period (otherwise the period of $A$ is bounded). If any unimodular system of equations over $A$ has a solution in $A$, then this is true for any quotient group of $A$. However, Proposition 1 shows that for $A_{p}$ this fact is not true. So, not all unimodular systems of equations over $A$ are solvable in $A$. One can even give an example of a system which has no solutions in $A$: if the projection of the system onto $A_{p}$ is the system from the proof of Proposition 1, then the system itself is unsolvable in $A$.
@@ 2.2 Positive case
Any $p$-nonsingular system of equations over an abelian group $A$ of prime period $p$ has a solution in $A$ itself.
@@
Look at the quotient group $A/pA$. Take the induced system $\{m_{i}=\overline{a}_{i}\}_{i\in I}$ of equations over $A/pA$, where $\overline{a}_{i}$ is the image of $a_{i}$ under taking the quotient $A\to A/pA$. It is a $p$-nonsingular system of equations over the abelian group $A/pA$ of period $p$, therefore, by Proposition 2 the system has a solution $\{\bar{c}_{j}\}_{j\in J}$ in $A/pA$.
@@
where $A_{p_{i}}$ is the $p_{i}$-component of $A$ (which has bounded period $p_{i}^{n_{i}}$). For the components $A_{p_{i}}$ the statement of the lemma follows from Proposition 3, therefore, it is also true for their direct sum. ∎
@@ 2.3 The final criterion
Now we can prove the criterion formulated in the introduction.
@@
Decompose $A$ into a direct sum of a divisible abelian group and a reduced one. The divisible abelian group has a solution of every nonsingular system of equations over itself (by Lemma 1). Therefore, nonsingular (particularly, unimodular) systems of equations have a solution in $A$ if and only if the same is true for its reduced part. Hence we can assume $A$ to be reduced.
@@ 3 Nilpotent groups
Now we switch back to the multiplicative notation. The results of Section 2 can be used to research solvability of systems of equations in nilpotent groups.
@@ 3.1 Periodic nilpotent groups
Take a look at the system $\{\overline{w}_{i}=1\}_{i\in I}$ over $G/Z(G)$, where for each $i\in I$ the word $\overline{w}_{i}$ is the word $w_{i}$ with coefficients changed to their images under taking the quotient $G\to G/Z(G)$. The system $\{\overline{w}_{i}=1\}$ is a unimodular system of equations over $G/Z(G)$ (which is a nilpotent group of class $l$ and of bounded period), therefore, the system is solvable in the quotient group $G/Z(G)$ itself by induction hypothesis.
This means that we can make a substitution of the variables in the system $\{w_{i}=1\}$ over $G$ such that for each $i\in I$ product $b_{i}$ of coefficients of the word $w_{i}$ lies in $Z(G)$ (i.e. the image of $b_{i}$ under taking the quotient $G\to G/Z(G)$ is $1$). So we can assume that each word $w_{i}$ has this property initially (that is, for every $w_{i}$, the product of the coefficients of $w_{i}$ lies in $Z(G)$).
@@ 3.2 Divisible nilpotent groups
Take a look at the system $\{\overline{w}_{i}=1\}_{i\in I}$ over $G/Z(G)$, where for each $i\in I$ the word $\overline{w}_{i}$ is the word $w_{i}$ with coefficients changed to their images under taking the quotient $G\to G/Z(G)$. The system $\{\overline{w}_{i}=1\}$ is a nonsingular system of equations over $G/Z(G)$ (which is a divisible nilpotent group of class $l$), therefore, the system is solvable in the quotient group $G/Z(G)$ itself by induction hypothesis.
This means that we can make a substitution of the variables in the system $\{w_{i}=1\}$ over $G$ such that for each $i\in I$ the product $b_{i}$ of coefficients of the word $w_{i}$ lies in $Z(G)$ (i.e. the image of $b_{i}$ under taking the quotient $G\to G/Z(G)$ is $1$). So we can assume that each word $w_{i}$ has this property initially (that is, for every $w_{i}$, the product of the coefficients of $w_{i}$ lies in $Z(G)$).
@@ 4 Open questions
The same question for mixed abelian groups, i.e. abelian groups which have both elements of infinite order and nontrivial elements of finite order.